# Mitigating the Linguistic Gap with Phonemic Representations for Robust Multilingual Language Understanding

### v1

@@ Abstract
Approaches to improving multilingual language understanding often require multiple languages during the training phase, rely on complicated training techniques, and—importantly—struggle with significant performance gaps between high-resource and low-resource languages. We hypothesize that the performance gaps between languages are affected by linguistic gaps between those languages and provide a novel solution for robust multilingual language modeling by employing phonemic representations (specifically, using phonemes as input tokens to LMs rather than subwords). We present quantitative evidence from three cross-lingual tasks that demonstrate the effectiveness of phonemic representation, which is further justified by a theoretical analysis of the cross-lingual performance gap [^1] Our code is available here: https://anonymous.4open.science/r/ipa-for-multilingual-nlu/README.md.
@@
for Robust Multilingual Language Understanding
Hae Ji Jung Korea University gpwl0709@korea.ac.kr                        Changdae Oh University of Seoul                        Jooeon Kang Sogang University                        Jimin Sohn GIST
Kyungwoo Song Yonsei University                        Jinkyu Kim Korea University                        David R. Mortensen Carnegie Mellon University dmortens@cs.cmu.edu
@@ 1 Introduction
In an era of large language models, natural language processing has promised to bring us together. However, there are gaps between how human language technologies perform in some languages (like English) and others (most of the languages of the world). Language technologies exhibit performance gaps between languages. To some degree, this is due to differences in resourcedness, but we hypothesize that linguistic gaps—the chasms that separate languages from one another structurally and lexically—also play a role. We further hypothesize that techniques that reduce linguistic gaps will also reduce performance gaps. In this paper, we focus on one such technique: representing languages phonetically using the International Phonetic Alphabet (IPA).
Figure 1 illustrates our motivation for using IPA as a universal language representation. The key contributions of this paper are as follows:
  • We demonstrate the advantage of using phonemic representation (i.e., IPA) for language modeling, particularly as a robust multilingual representation.
@@
  • We empirically validate the effectiveness of phonemic representations by comparing the performance gap and linguistic gap across languages with subword or character representations.
  • We further explain the empirical observations with theoretical analysis from domain generalization literature, by treating the linguistic gap as the domain gap caused by lexical and syntactic disparities.
Figure 1: Example of word, character, and phoneme units for a sentence (English and Korean).
2 Related Work
Cross-lingual Transfer Learning
approaches aim to transfer knowledge from one language to enhance NLP performance in another. While adversarial training methods Lample et al. (2018); Chen et al. (2018, 2021); Yu et al. (2023) attempt to learn a language-agnostic representation, their training can be unstable Balcan et al. (2023). Non-adversarial joint learning approaches Cotterell and Heigold (2017); Gu et al. (2018); Wei et al. (2021); Zheng et al. (2021) also present promising results, but necessitate significant multilingual data. Recently, Yang et al. (2022) revealed the empirical correlation between cross-lingual transferability and representation discrepancy, yet they do not provide any theoretical statement to justify such numerical analysis, which is presented in this work.
Multilingual Language Models (MLLMs)
have been rapidly advanced with a popularized large-scale pre-training and fine-tuning paradigm Devlin et al. (2018); Conneau and Lample (2019); Conneau et al. (2020); Chi et al. (2021); Workshop et al. (2022); Wei et al. (2023). Although MLLMs have demonstrated remarkable potential in diverse multilingual tasks, there is a significant performance gap between different languages, especially in the case of high-resource language versus low-resource language Wu and Dredze (2020); Zhao et al. (2023). In this work, we propose a novel paradigm for robust multilingual language modeling: IPA as a universal language representation.
Phonemic Language Representation
has been applied in previous studies Bharadwaj et al. (2016); Chaudhary et al. (2018); Dalmia et al. (2019); Hu et al. (2019); Nguyen et al. (2023), but primarily for speech recognition tasks or where they were not focused on the linguistic gap between languages. Our work revisits phonemic representation for MLLMs to mitigate the linguistic gap for the first time.
@@ 3 Experimental Setup
In this section, we briefly describe the experiment setup in terms of models, datasets, and tasks. Refer to Appendix A for further details.
3.1 Models and Data Preprocessing
@@
We compare LMs with three different types of language representation: subword, character, and phoneme. All models222pre-trained weights were obtained from https://huggingface.co/models are pre-trained on multilingual data that covers around 100 languages from Wikipedia dump files. We employ an off-the-shelf MLLMs, multilingual BERT (mBERT; Devlin et al. (2018)). For a subword-based model which is trained with byte-pair encoding. For a character-based model, we utilize CANINE Clark et al. (2022), which is a tokenization-free LM that directly maps each character to its codepoint by hashing. It is pre-trained on the same data and training objectives as mBERT. For a phoneme-based model, we adopt XPhoneBERT Nguyen et al. (2023), which has the same model architecture as mBERT.
@@
While character-level models are known to better generalize to low-resource languages Clark et al. (2022), their general performance falls behind subword-based models. For a fair comparison between the representations, we primarily compare phoneme-based model to character-based one instead of directly comparing it to the subword-based model (i.e., mBERT), and leave further improvements of overall performance as future work.
@@
In order to prepare inputs for a phoneme-based model, we employed G2P (Grapheme-to-Phoneme) conversion to obtain an IPA version of the input. This conversion was done with Epitran [^3] https://github.com/dmort27/epitran Mortensen et al. (2018), an external tool for G2P conversion. After converting to IPA, we inserted white space between every character to make it compatible with XPhoneBERT’s tokenizer.
3.2 Downstream Tasks
We adopted the cross-lingual generalization benchmark tasks suggested in XTREME Hu et al. (2020). Among them, we selected two types of tasks; classification and structured prediction. For evaluation languages, we choose eight languages of three groups—high-resource (eng, deu, fra), low-resource (urd, hin, swa), and typologically distinct (kor, ukr)—to analyze the impact of phonemic representation with respect to each category.
Sentence-level Classification.
We employ XNLI Conneau et al. (2018) dataset, which is a representative benchmark for the natural language inference task on the cross-lingual generalization setting.
Token-level Classification.
We choose POS tagging and NER as our testbed for structured prediction tasks, both requiring predicting labels for each word in sentences. We utilize the corpora from Universal Dependencies [^4] https://universaldependencies.org/ , v2.13, 148 languages, released Nov 15, 2023. for POS tagging, and WikiAnn Pan et al. (2017) with train, dev, test splits following Rahimi et al. (2019) for NER.
4 Results
@@
($\Delta$ from ENG) ($\Delta$ from ENG)
Subword 80.80 62.93 (17.87) 61.57 (19.23)
Character 75.02 59.72 (15.30) 56.55 (18.46)
Phoneme 71.89 60.88 (11.00) 56.10 (15.78)
@@
Table 1: Accuracy (%) on XNLI task. ENG, SWA, URD refer to English, Swahili, and Urdu, respectively. Phonemic representation shows relatively small performance gaps compared to other representations.
Method Language Performance gap Linguistic Gap
EN DE FR KO UK UR HI SW Std. ($\downarrow$) Mean diff. ($\downarrow$) Mean CKA ($\uparrow$)
Named Entity Recognition
Character 88.80 77.99 90.87 80.17 85.61 75.88 73.33 74.84 6.68 0.10 0.34
Phoneme 91.36 81.36 91.93 79.33 83.67 76.41 82.19 78.06 5.79 0.08 0.43
Part-of-Speech Tagging
Character 73.54 81.20 80.55 70.1 84.28 93.15 - - 8.14 0.12 0.30
Phoneme 70.50 78.68 78.10 77.35 83.75 93.96 - - 7.88 0.12 0.52
Table 2: Performance of POS tagging and NER across different languages. Std. refers to the standard deviation of the scores across the languages, and Mean diff. indicates average pairwise difference of F1 scores. Mean CKA represents the average linguistic gap between languages.
@@
Figure 2: Linguistic gaps across languages. Upper and lower triangular elements indicate pairwise linguistic gaps derived with phoneme-based model and character-based model, respectively. Lighter color indicates larger CKA score, which means smaller discrepancy. Upper triangular elements show relatively lighter colors, implying smaller discrepancies across languages.
@@
Figure 3: Qualitative analysis of performance gap (difference of accruacy) on XNLI task. (Left) the absolute difference between performance across two languages, (center) centered kernel alignment (CKA) scores to measure cross-lingual embedding similarity, and (right) Sinkhorn distance on the output probability space. Phonemic representation shows relatively small performance gaps w.r.t. EN $\leftrightarrow$ SW and EN $\leftrightarrow$ UR, and these gaps are correlated with similarity and discrepancy on the embedding space (CKA) and logit space (Sinkhorn distance).
Phoneme-based Model on Low-Resource Languages.
The phoneme-based model shows promising results compared to other models, especially for low-resource languages. As shown in Table 1, the phoneme-based model has the smallest performance gap between English and other low-resource languages – swa, urd. Furthermore, while subword-based mBERT achieves the highest scores, the performance disparity across models narrows when it comes to low-resource languages. Table 2 also suggests that the phoneme-based model exhibits superiority in addressing low-resource languages. For NER, on languages like urd, hin, and swa, the phoneme-based model significantly outperforms character-based model, highlighting the capability of the phoneme-based model’s generalization over the low-resource languages.
Performance Gap Across Languages.
We observe that the phoneme-based model performs the most consistently across languages. The leftmost panel in Figure 3, shows how each language representation results in performance gaps across different languages. Here, the phoneme-based model comes with the lowest performance gap. Table 2 also shows the phoneme-based model’s robustness in terms of performance gap across languages. From the table, standard deviation (Std.) and mean difference (Mean diff.) both indicate how the results for all languages differ from each other. The phoneme-based model ends up with a smaller standard deviation in both NER and POS tasks.
Linguistic Gap of Different Representations.
To analyze the potential of phonemes as a robust representation for multilingual language modeling, we check the gap between different languages over different representations. As in Yang et al. (2022), we use centered kernel alignment (CKA) to compute the representation similarity. Figure 3 shows that phonemic representation shows higher CKA and lower Sinkhorn distance compared to other representations, meaning that the phonemic representations from different languages are relatively closer to each other than those of subword or character representations.
Figure 2 also illustrates the linguistic gap between languages with their pairwise similarity. After fine-tuning each model on downstream tasks, we compute the discrepancy between different languages using held-out parallel data. It can be clearly stated that the upper triangle part (which corresponds to phonemic representations) of the map has large values that imply a smaller linguistic gap, contributing to a smaller performance gap.

### v3

# Mitigating the Linguistic Gap with Phonemic Representations for Robust Cross-lingual Transfer

@@ Abstract
Approaches to improving multilingual language understanding often struggle with significant performance gaps between high-resource and low-resource languages. While there are efforts to align the languages in a single latent space to mitigate such gaps, how different input-level representations influence such gaps has not been investigated, particularly with phonemic inputs. We hypothesize that the performance gaps are affected by representation discrepancies between these languages, and revisit the use of phonemic representations as a means to mitigate these discrepancies. To demonstrate the effectiveness of phonemic representations, we present experiments on three representative cross-lingual tasks on 12 languages in total. The results show that phonemic representations exhibit higher similarities between languages compared to orthographic representations, and it consistently outperforms grapheme-based baseline model on languages that are relatively low-resourced. We present quantitative evidence from three cross-lingual tasks that demonstrate the effectiveness of phonemic representations, and it is further justified by a theoretical analysis of the cross-lingual performance gap.
@@
for Robust Cross-lingual Transfer
Haeji Jung1, Changdae Oh2, Jooeon Kang3, Jimin Sohn4, Kyungwoo Song5, Jinkyu Kim1, David R. Mortensen6 1Korea University, 2University of Wisconsin–Madison, 3Sogang University, 4GIST, 5Yonsei University, 6Carnegie Mellon University E-mail: gpwl0709@korea.ac.kr, dmortens@cs.cmu.edu
@@ 1 Introduction
Large language models have significantly advanced natural language processing, offering improved capabilities across numerous languages. However, substantial performance gaps remain, particularly between high-resource languages like English and the majority of the world’s low-resource languages. While these gaps are partly driven by discrepancies in data availability and quality, recent studies suggest that linguistic gaps—potentially caused by structural and lexical differences—also contribute significantly to these disparities .
Cross-lingual transfer techniques, which aim to adapt to arbitrary target language, have shown promise with the advancement of pre-trained multilingual language models Devlin et al. (2019); Conneau et al. (2020); Clark et al. (2022). However, they continue to face challenges, particularly with low-resource languages. One line of prior research has focused on mitigating these gaps through cross-lingual representation alignment Zhang et al. (2022); Wu and Monz (2023); Stap et al. (2023), but these efforts often overlook the impact of varying input representations on performance consistency across languages.
In this work, we explore the use of phonemic representations written in International Phonetic Alphabet (IPA) characters as a robust input representation (see Figure 1) to reduce linguistic gaps and, consequently, performance gaps across languages. We define the linguistic gap as the representation discrepancy between embedding vectors and the performance gap as the relative difference in downstream task performances between languages, to analyze the impact of phonemic representations in cross-lingual adaptation.
Figure 1: Example of orthographic and phonemic input representations of a sentence (English and Korean).
Our empirical analysis shows that phonemic representations consistently reduce linguistic gaps between languages compared to orthographic character-based models. This reduction in linguistic gaps directly correlates with smaller performance gaps in tasks such as cross-lingual natural language inference (XNLI), named-entity recognition (NER), and part-of-speech (POS) tagging, demonstrating the potential of phoneme-based models to enhance cross-lingual transfer across diverse languages. We further support these findings with theoretical analysis from domain generalization literature, where we frame the performance gap as a consequence of linguistic gaps driven by lexical and syntactic differences.
Our key contributions are as follows:
  • We revisit the use of phonemic representations (IPA) as a universal input strategy to reduce performance gaps across languages in multilingual language models.
@@
  • We empirically demonstrate the effectiveness of phonemic representations by comparing them with subword and character-based models, highlighting their ability to minimize both performance and linguistic gaps.
  • We provide a theoretical explanation for the observed benefits of phonemic representations, drawing parallels between linguistic gaps in multilingual settings and domain gaps in domain generalization literature.
2 Related Works
2.1 Cross-lingual Transfer with Multilingual Language Model
Cross-lingual transfer learning aims to improve performance on low-resource languages (LRLs) by leveraging data from high-resource languages. Models like mBERT Devlin et al. (2019) and XLM-R Conneau et al. (2020), trained on hundreds of languages, have demonstrated effective cross-lingual adaptation by leveraging large multilingual pre-train datasets Fujinuma et al. (2022); Wu and Dredze (2020); Conneau et al. (2020). However, significant performance discrepancies remain between languages due to differences in data availability, script types, and language families Wu and Dredze (2020); Muller et al. (2021a); Bagheri Nezhad and Agrawal (2024). This "performance gap" has been systematically evaluated in benchmarks such as XTREME Hu et al. (2020), highlighting the need for methods that can ensure more consistent performance across languages.
2.2 Cross-lingual Representation Gap
One approach to reducing performance gaps focuses on narrowing the representation gap between languages. Multilingual pre-training enables models to learn shared representation space for multiple languages. Singh et al. (2019) and Muller et al. (2021b) both analyze the representations of pre-trained multilingual models and observe that lower layers are responsible for this cross-lingual alignment. Yang et al. (2022) employs mixup Zhang et al. (2018) to bring representations closer together, improving performance by reducing their distance in the latent space. Other works show a strong correlation between representation distance and machine translation performance, suggesting that improved alignment leads to better transfer results Wu and Monz (2023); Stap et al. (2023). While these studies provide valuable insights into the benefits of aligning cross-lingual representations, they do not explore how variations in input-level representations, such as the use of phonemic representations instead of orthographic characters, might affect this alignment. This paper investigates how phonemic representations can further reduce cross-lingual gaps.
2.3 Phonemic Representations for Multilingual Language Modeling
Phonemes, typically represented by International Phonetic Alphabet (IPA) characters, are the perceptual sounds of a language. Phonemic representations offer a language-agnostic input that can enhance multilingual modeling, especially for LRLs. By using phonological features that are less dependent on specific orthographic systems, these representations offer a language-agnostic alternative that can help bridge performance gaps across languages. Previous studies have shown that using the IPA characters as input can enhance performance in cross-lingual tasks such as named entity recognition Chaudhary et al. (2018); Bharadwaj et al. (2016); Leong and Whitenack (2022) and machine translation Chaudhary et al. (2018); Sun et al. (2022), particularly for low-resource languages. Similarly, Sohn et al. (2024) report that phoneme-based models outperform other baselines on target languages unseen during pre-training. While these works demonstrate the potential of phonemic representations in language modeling, few have explored the specific embeddings and representations of phonemes. Although some studies have developed pre-defined phoneme embeddings (e.g., PanPhon Mortensen et al. (2016), Phoible Moran and McCloy (2019)) and learned embeddings from masked language modeling Li et al. (2023); Jia et al. (2021); Sundararaman et al. (2021); Zhang et al. (2022), there is limited understanding of how these embeddings function in cross-lingual contexts.
We utilize XPhoneBERT Nguyen et al. (2023), a model pre-trained with phonemes across approximately 100 languages, to investigate how using phonemic representations as input can mitigate cross-lingual performance discrepancies. Our empirical and theoretical analyses provide new insights into the benefits of phonemic representations for multilingual language modeling, particularly in terms of narrowing the cross-lingual linguistic gap and performance gap.
@@ 3 Experimental Setup
In this section, we describe the experiment setup in terms of models, datasets, and downstream tasks, including the selected target languages and details for preprocessing. Additionally, we provide details on evaluation strategies, particularly on quantifying the performance and linguistic gap.
3.1 Models
@@
We employ three masked language models that are pre-trained on multilingual corpus that covers around 100 languages from Wikipedia dump files111pre-trained weights are obtained from https://huggingface.co/models: mBERT Devlin et al. (2019), CANINE Clark et al. (2022), and XPhoneBERT Nguyen et al. (2023). Each model is trained on different types of language representation.
Multilingual BERT (mBERT) is a subword-based model that utilizes WordPiece algorithm for tokenization. During pre-training, mBERT learns to perform masked language modeling (MLM) and next sentence prediction (NSP).
CANINE is a multilingual character-based model that is trained on the same corpus with the same training objective as mBERT. CANINE is a tokenization-free language model that directly maps each unicode character to its codepoint by hashing. This prevents unknown tokens, enabling the model to handle a large amount of distinct characters.
Lastly, XPhoneBERT is a phoneme-based model trained to do MLM. XPhoneBERT follows the pre-training scheme of XLM-R Conneau et al. (2020), so NSP is not employed in its pre-training. This model takes as input the sequence of IPA characters, where the input data are created from original text by G2P conversion followed by phoneme segmentation.
@@
While character-level models are known to better generalize to low-resource languages Clark et al. (2022), their general performance falls behind subword-based models. To specifically compare input representations–phonemes versus orthographic scripts–we minimize the impact of different tokenization units by focusing on phoneme-based models versus character-based models, rather than directly comparing with subword-based models like mBERT. Nevertheless, we include mBERT results for the XNLI task to highlight its significant performance drop on low-resource languages. For other tasks, we report results from phoneme and character-based models to ensure a fair comparison, and leave further improvements of character-level models in overall performance as future work.
3.2 Downstream Tasks
We adopted the cross-lingual generalization benchmark tasks suggested in XTREME Hu et al. (2020).
Token-level Classification.
We choose POS tagging and NER as our testbed for structured prediction tasks. Both tasks require labeling each token from the model. These types of tasks were previously analyzed as being relatively independent from the data size of each language used for pre-training the language model Hu et al. (2020). We find this particularly suitable in our scenario where two models with different pre-training strategy are compared. For datasets, we utilize the corpora from Universal Dependencies [^2] https://universaldependencies.org/ , v2.13, 148 languages, released Nov 15, 2023. for POS tagging, and WikiAnn Pan et al. (2017) with train, dev, test splits following Rahimi et al. (2019) for NER.
Sentence-level Classification.
XTREME supports two sentence-level classification tasks. This type of task requires semantic understanding of given sentences to make a prediction. We employ XNLI Conneau et al. (2018) dataset, which is a representative benchmark for the natural language inference task on cross-lingual generalization setting. This task requires the model to classify the relation of two given sentences into three different classes.
3.3 Performance Gap
We analyze performance gaps of each model for all downstream tasks. As we are interested in how different models with different input types performs consistent across languages rather than their absolute overall performance, we take the relative percentage difference (RPD) Miller (2011) to derive the performance gap. Here, we define RPD as
$\text{RPD}(L_{i},L_{j})=\frac{|\text{S}(L_{i})-\text{S}(L_{j})|}{\frac{1}{2}(\text{S}(L_{i})+\text{S}(L_{j}))}\times 100,$ (1)
where S$(L_{i})$ represents the performance for the language $L_{i}$. This is used to analyze the performance gap, which specifically computes the relative performance gaps across languages.
3.4 Linguistic Gap
To compute representation discrepancy across languages, we use FLORES+ Costa-jussà et al. (2022) corpus which contains parallel sentences of more than 200 languages. We employ devtest set of each language subset, which contains 1,012 sentences.
After training each model on each downstream task, we utilize each model to obtain similarity in their representations. We adopt mean-pooling to obtain sentence representations and Centered Kernel Alignment (CKA) Kornblith et al. (2019) to measure the similarity, which Del and Fishel (2021) has recommended for robust analysis on cross-lingual similarity. CKA is defined as,
$\text{CKA}(\mathbf{X},\mathbf{Y})=\frac{\|\mathbf{X}^{T}\mathbf{Y}\|_{2}^{2}}{\|\mathbf{X}^{T}\mathbf{X}\|_{2}\|\mathbf{Y}^{T}\mathbf{Y}\|_{2}},$ (2)
where features $\mathbf{X}$ and $\mathbf{Y}$ are from different languages. They are extracted from the input embedding layers as we are interested in how different input types (i.e., orthographic vs. phonemic) affect cross-lingual alignment, and Muller et al. (2021b) finds that cross-lingual alignment happens in the lower layers of the model. We use this similarity scores computed with CKA to refer to linguistic gaps, where smaller CKA score means larger linguistic gap.
3.5 Implementation Details
Models were trained for 30 epochs on a single NVIDIA A5000 GPU for POS tagging, 30 epochs a single NVIDIA A40 GPU for NER, and 20 epochs on NVIDIA A6000 for XNLI. For all experiments, batch size was set to 128 and AdamW Loshchilov and Hutter (2018) optimization was used. Additionally, cosine learning rate scheduler was adopted with its initial learning rate set by grid search. Learning rates used for each model on each language are in the supplementary material.
3.6 Data Preparation
Languages.
To evaluate token-level tasks, we selected 10 languages with diverse typoloigical background—English(eng), French(fra), Russian(rus), Italian(ita), Hungarian(hun), Ukrainian(ukr), Korean(kor), Turkish(tur), Finnish(fin), and Hindi(hin). First four languages are high-resource languages, where English, French, and Italian are written in Latin scripts and Russian in Cyrillic. The other languages are pre-trained on each model with moderate or small amount of data, and are written in diverse scripts, such as Hangul, Cyrillic and Devanagari. For further analysis using sentence-level tasks, we chose two low-resource languages—Swahili(swa), and Urdu(urd)—to compare with a representative high-resource language, English(eng).
@@ 3.6 Data Preparation
In order to prepare inputs for a phoneme-based model, we employed G2P (Grapheme-to-Phoneme) conversion to obtain an IPA version of the input. This conversion was done with Epitran [^3] https://github.com/dmort27/epitran Mortensen et al. (2018) , an external tool for G2P conversion. After converting to IPA, phoneme segmentation with a python package, segments [^4] https://pypi.org/project/segments/, to identify each phoneme. Lastly, to make it compatible with XPhoneBERT’s tokenizer, white space was inserted between every phoneme.
4 Results and Analysis
Here, we present our observations and analyses of the results. We first discuss the behavior of phoneme-based model towards low-resource languages and writing systems, which contributes to robust cross-lingual performance. Next, we delve into the performance and linguistic gaps of phoneme-based models through empirical and theoretical analyses.
Method Language Performance gap Linguistic Gap
eng fra rus ita hun ukr kor tur fin hin Std. ($\downarrow$) Mean RPD ($\downarrow$) Mean CKA ($\uparrow$)
Named Entity Recognition
Character 87.13 91.27 91.80 92.26 93.14 93.88 84.11 92.92 90.45 87.68 0.0316 4.02 0.4584
Phoneme 83.61 89.42 89.60 90.56 91.89 92.76 87.19 92.35 89.23 88.23 0.0259 3.52 0.7195
Part-of-Speech Tagging
Character 96.62 95.54 87.91 96.06 74.57 85.79 86.71 90.49 91.78 96.81 0.0692 8.77 0.4593
Phoneme 95.94 96.35 86.69 96.37 85.87 91.32 85.82 91.11 93.76 96.94 0.0455 5.80 0.7204
Table 1: Performance of POS tagging and NER across different languages. Std. refers to the standard deviation of the scores across the languages, and Mean RPD indicates average relative difference of F1 scores between different languages. Mean CKA represents the average linguistic gap between languages.
4.1 Phoneme-based Model on Low-Resource Languages and Writing Systems
We observe that phoneme-based model shows promising performance in low-resource languages and writing systems (scripts). Results from Table 1 show that phoneme-based model outperforms the character-based model on NER task, in languages written in scripts other than major scripts [^5] Latin and Cyrillic are scripts that are used the most during the pre-training phase.—Korean and Hindi. This can be attributed to the fact that named entities, such as geopolitical or personal names, are often pronounced similarly across languages. When different writing systems and scripts are used, models may struggle to align such entities. However, representing them in IPA characters that reflect their pronunciations helps the model to better align these entities, resulting in better cross-lingual transfer. This results align with findings from Muller et al. (2021a); Sohn et al. (2024), which focus on unseen languages, whereas we observe this phenomenon with diverse ‘seen’ languages.
Results also demonstrates the potential of phoneme-based model in addressing low-resource languages. As shown in Table 2, the phoneme-based model achieves a smaller gap when transferred to low-resource languages such as Swahili and Urdu, compared to other baselines. This finding is further analyzed in Section 4.2
4.2 Performance Gap Across Languages
@@
Acc. Acc. $\Delta$ from eng Acc. $\Delta$ from eng
(Rel./Abs.) (Rel./Abs.)
Subword 80.80 62.93 24.87 / 17.87 61.57 27.01 / 19.23
Character 75.02 59.72 22.71 / 15.30 56.55 28.08 / 18.47
Phoneme 71.89 60.88 16.59 / 11.01 56.10 24.67 / 15.79
@@
Table 2: Accuracy (%) and relative/absolute performance gaps on XNLI task. eng, swa, and urd refer to English, Swahili, and Urdu, respectively, and relative difference is computed with RPD. Phonemic representation shows relatively small performance gaps compared to other representations.
@@
Figure 2: Linguistic gaps across languages in each model. (Center) Upper and lower triangular elements of the heatmap indicate pairwise linguistic gaps derived with character-based model and phoneme-based model, respectively. Darker color indicates larger CKA score, which means smaller discrepancy. Lower triangular elements show relatively darker colors, implying smaller discrepancies across languages of phoneme-based model. (Left, right) T-SNE plots for each model are shown with only five languages, for better visibility.
We observe that the phoneme-based model consistently exhibits the smallest performance gap across diverse languages, highlighting its robustness in cross-lingual tasks. In Table 1, we present the standard deviation (Std.) and average percentage difference (Mean diff.) for all models, which reflect the variability in performance across different languages. The phoneme-based model exhibits both a lower standard deviation and a smaller average percentage difference in the NER and POS tasks, demonstrating its relatively stable performance across different languages.
Table 2 provides additional evidence by showing that the phoneme-based model achieves a smaller gap in performance between English and other low-resource languages—Swahili(swa) and Urdu(urd)—compared to other models. We report both relative and absolute differences in performance, with the relative difference calculated as described in Section 3.3.
While subword-based mBERT achieves the highest scores, the performance gaps between models narrow when applied to low-resource languages, with outperforming the phoneme-based model by 8.91% in English and by 2.05% and 5.47% in Swahili and Urdu, respectively. This reflects subword LM’s significant performance drops on low-resource languages, while highlighting the phoneme-based LM’s robustness in cross-lingual transfer to such languages. The leftmost panel of Figure 3 also illustrates the performance gaps of each model, where the phoneme-based model predominantly displays lower gaps compared to others.
These metrics collectively suggest that phonemic representations offer a more consistent performance in multilingual settings, reducing the disparities typically observed when models are applied to languages with varying resource availability.
@@
Figure 3: Qualitative analysis of performance gap (difference of accruacy) on XNLI task. (Left) the absolute difference between performance across two languages, (center) centered kernel alignment (CKA) scores to measure cross-lingual embedding similarity, and (right) Sinkhorn distance on the output probability space. Phonemic representation shows relatively small performance gaps w.r.t. eng $\leftrightarrow$ swa and eng $\leftrightarrow$ urd, and these gaps are correlated with similarity and discrepancy on the embedding space (CKA) and logit space (Sinkhorn distance).
4.3 Linguistic Gap of Different Representations
To investigate the potential of phonemes as a robust representation for multilingual language modeling, we analyze the linguistic gap between languages using different input representations. Following Yang et al. (2022); Muller et al. (2021b), we use linear CKA to quantify representation similarity across languages. Figure 2 shows the pairwise similarities between languages, with the lower triangle of the heatmap, which corresponds to phonemic representations, demonstrating higher similarity values. This indicates a smaller linguistic gap compared to models that use orthographic inputs, contributing to a smaller performance gap. Moreover, the t-SNE plots placed in both sides show how the distributions of the representations from different languages resemble each other. Phoneme-based model exhibits more similar distribution across languages.
Figure 3 further supports these observations by showing the linguistic gap after fine-tuning on the XNLI task. The plot in the center illustrates that phonemic representations have higher CKA scores than other baseline models, indicating closer alignment between language representations. As XNLI directly learns to build a sentence representation during fine-tuning, we extract the representation from the last hidden layer unlike in other token-level tasks. Additionally, by using Sinkhorn distance to compare the logit space, we observe that the phoneme-based model shows lower distances, reflecting more consistent predictions across languages.
These results highlight the potential of phonemic representations to address the performance gaps that challenge multilingual language models, particularly in bridging the gap between high-resource and low-resource languages by more similar representations.
4.4 Connecting Performance Gap and Linguistic Gap
Correlation Spearman’s R Kendall’s T
coefficient p-value coefficient p-value
Performance Gap <-> eng Performance 0.111 5.60E-01 0.104 4.30E-01
Performance Gap <-> S-Dist 0.681 3.50E-05 0.457 2.00E-04
Performance Gap <-> CKA -0.782 3.40E-07 -0.577 2.10E-06
Table 3: Correlation analysis with 45 phoneme-based models. We fine-tune the phoneme-based language model XPhoneBERT on three languages, eng, swa, and urd, with 15 different random seeds and conduct two types of correlation analyses.
Correlation Analysis.
Meanwhile, one may speculate the low-performance gap of the phoneme-based model can be driven by the low English performance rather than reducing the linguistic gap. To clarify this, we simulate 15 repeated runs (with different random seeds) of phonemic representation using 10% of the XNLI train dataset over English, Swahili, and Urdu. After computing the best performance per each language, Sinkhorn distance (S-Dist), and CKA between English and the other two languages, we conducted correlation analyses by performing hypothesis tests with Spearman’s rank correlation coefficient and Kendall’s Tau.
As can be seen from Table 3, rather than the English performance, S-Dist and CKA have stronger correlations, indicating that the linguistic gap has stronger correlations that are statistically significant (with a significant level less than 0.01).